OpenStudy (anonymous):

Last study guide question! I have sorted worked through some it but want to make sure... How many grams of C2H5OH are used up when 6.25 grams of O2 react according to the same equation? C2H5OH + 3O2 = 2CO2+ 3H2O

OpenStudy (anonymous):

I know the weight is 276.408

OpenStudy (anonymous):

and the molar mass is 46.068

OpenStudy (anonymous):

the moles are 6. If you could just share the equation, I want to work through it by myself.

thomaster (thomaster):

Convert 6.25 gram O2 to moles, check the mole ratio calculate the amount of moles C2H5OH that is needed to react convert that to gram :)

OpenStudy (anonymous):

i multiply 6.25 by 6.023x10^23 to get moles correct?

thomaster (thomaster):

Mass to mole

thomaster (thomaster):

You got the amount of moles \(\sf\large O_2\)?

OpenStudy (anonymous):

0.39064941558847 i dont think thats right lol

OpenStudy (anonymous):

because the molar mass of oxygen is 15.999

thomaster (thomaster):

the molecular mass of \(\sf O_2\) is 31.998 g/mol (15.999*2) So \(\sf\large\dfrac{6.25}{31.998}=0.1953 ~mol\)

OpenStudy (anonymous):

oh because you have 02 so 15.999x2=31.998

thomaster (thomaster):

Yes :) Now we have to calculate how many moles of ethanol is used. 

OpenStudy (anonymous):

ahhh the cross table :) 0.1953x1/3= 0.0651

OpenStudy (anonymous):

calculating those moles to weight we get =2.999055444

OpenStudy (anonymous):

then subtracting 6.25-2.999055444= 3.25094456?

thomaster (thomaster):

yea 0.0651*46.07 =2.999 :) great job

thomaster (thomaster):

gram sorry

OpenStudy (anonymous):

oh 2.999 is my answer?

thomaster (thomaster):

yes :)

thomaster (thomaster):

2.999 gram ethanol reacts with 6.25 gram oxygen

OpenStudy (anonymous):

why do i always think these problems are much harder then they really are hahaha

thomaster (thomaster):

Yea when you practice some of these questions, you'll realize that it's actually very easy :) Save the mol conversions diagram, it can be very helpful for these questions.

OpenStudy (anonymous):

i saved it to my desktop. it really clears it up. Thanks a lot once again!